OpenStudy (anonymous):

When uranium-238 emits an alpha particle, what isotope is created? Show the nuclear equation for this reaction.

OpenStudy (joannablackwelder):

Do you know what an alpha particle is?

OpenStudy (michele_laino):

we have the subsequent nuclear decay: \[\Large {U^{238}} \to T{h^{234}} + {\alpha ^4},\quad 4.5 \times {10^9}years\]

OpenStudy (anonymous):

i have no idea what an alpha particle is

OpenStudy (joannablackwelder):

http://www.chemteam.info/Radioactivity/Writing-Alpha-Beta.html

OpenStudy (michele_laino):

an alpha particel is the nucleus of the deuterium, which contains 2 protons and 2 neutrons, so its mass number is 2+2=4, furthermore, since an alpha particle doesn't have electrons, it charge is +4 e, where e=1.6*10^(-19) Coulombs

OpenStudy (michele_laino):

particle*

OpenStudy (joannablackwelder):

Alpha particles are typically written as  And yes, 2 protons and 2 neutrons leave the nucleus of the atom undergoing alpha decay.

OpenStudy (anonymous):

ok so i would add that to uranium

OpenStudy (joannablackwelder):

No, as @Michele_Laino showed, the alpha particle is one of the products.

OpenStudy (joannablackwelder):

And we can figure out with the other product is by the number of protons left.

OpenStudy (michele_laino):

you have to add the alpha particle to Thorium 234

OpenStudy (joannablackwelder):

Since the uranium is losing 2 protons (it started with 92), it has 90 left over.

OpenStudy (joannablackwelder):

Which is Th

OpenStudy (anonymous):

so its thorium 90 or am i missing another step

OpenStudy (joannablackwelder):

And it lost 4 from it's mass number (protons + neutrons), which is why the mass number of Th is 234.

OpenStudy (joannablackwelder):

It is Thorium-234

OpenStudy (anonymous):

Thank you!!

OpenStudy (michele_laino):

Thank you! :)

OpenStudy (joannablackwelder):

You're welcome. :-)

OpenStudy (joannablackwelder):

And, @Michele_Laino Please don't just give answers. That is against policy and does not help the student to learn most effectively.

OpenStudy (michele_laino):

I don't gave the direct answer, I also wrote the created isotope, as requested by the question

OpenStudy (michele_laino):

@JoannaBlackwelder

OpenStudy (michele_laino):

the unique way to know the requested isotope, is to look at the decays diagram, as I have done @JoannaBlackwelder

OpenStudy (jtfan2000):

This entire site just gives out answers. Should be called "do my homework for me"